# Learning nonnegative matrix factorizations from compressed data

### v1

@@ 3.4. Nonnegativity in compression
In the next section, we discuss the approaches to solve compressed nonnegative matrix factorization problems. In particular, we consider the variations of multiplicative updates algorithm to iteratively minimize the objective functions that we have formulated in this section. The multiplicative updates algorithm is valid due to the fact that the matrices involved in the iterative process are nonnegative. This convenient property is destroyed by sketching unless we have that ${\mathbf{A}}^{T}{\mathbf{A}}$ is an element-wise nonnegative matrix. While this is not expected to be true neither for approximately orthonormal random sketches nor for the data-adapted sketching matrices coming from randomized rangefinder algorithm, to overcome this issue, it suffices to add some extra penalty terms taking the form
$\sigma\|\mathbf{1}_{m}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}\quad\text{and/or}\quad\sigma\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\mathbf{1}_{n}\|^{2},$ (3.14)
@@
Corollary 3.12.
Suppose ${\mathbf{X}}$ has an exact nonnegative factorization ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$, where ${\mathbf{U}}_{0}\in{\mathbb{R}}^{n\times k}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}^{m\times k}$ and they are both full-rank, $k\leq\min\{n,m\}$. Let ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ are generic random matrices of sizes $k\times n$ and $m\times k$, respectively. If for some $\lambda_{1},\lambda_{2}>0$ and $\sigma_{1},\sigma_{2}\geq 0$
$\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}L$ $({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})$ (3.15)
@@
where $L({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}):=\|{\mathbf{A}}_{1}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}+\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{A}}_{2}\|_{F}^{2}+\lambda_{1}\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{A}}_{1}{\mathbf{X}}}^{\perp}\|_{F}^{2},$ then ${\mathbf{X}}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}$.
@@
and similarly for adding the term $\sigma_{2}\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\mathbf{1}_{n}\|^{2}_{F}$. Then the statement follows directly from Theorem 3.1. ∎
When we do not assume that ${\mathbf{X}}$ has exactly nonnegative rank $k$, adding the regularizer of the form (3.14) is still possible under an additional condition, essentially imposing that the column-sums of ${\mathbf{X}}$ and ${\mathbf{U}}{\mathbf{V}}^{T}$ approximately match if ${\mathbf{U}},{\mathbf{V}}$ are optimal to (1.1).
Corollary 3.13.
Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be a nonnegative matrix and ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ is a matrix with orthogonal rows. Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1.1) of rank $r$ and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. Additionally assume that $\left\|\mathbf{1}^{T}({\mathbf{X}}_{0}-{\mathbf{X}})/\mathbf{1}^{T}{\mathbf{X}}\right\|\leq\varepsilon<1$, where /// denotes element-wise division and $\mathbf{1}\in{\mathbb{R}}^{m}$ is the vector of all ones. If $\lambda,\sigma>0,$
$\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}=\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left\|{\mathbf{A}}\left({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\right)\right\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2},$ (3.16)
@@
$\frac{\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c_{\lambda}\left[\frac{\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|{\mathbf{X}}-P_{\mathbf{A}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\varepsilon^{2}\right].$
For example, if we have that $\lambda,\varepsilon\in(0,1/2)$ one can bound $c=\max\{6,8/\lambda\}.$
@@
Let ${\mathbf{D}}\in{\mathbb{R}}^{n\times n}$ be the diagonal matrix with nonzero entries given by $(\mathbf{1}^{T}{\mathbf{X}}/\mathbf{1}^{T}{\mathbf{X}}_{0})$, and so $\mathbf{1}^{T}{\mathbf{X}}_{0}{\mathbf{D}}=\mathbf{1}^{T}{\mathbf{X}}$. Note that ${\mathbf{X}}_{0}{\mathbf{D}}$ is some (not necessarily optimal) solution to (3.16) and so we have
$\|P_{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}$ $\leq\|P_{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}$
$\leq\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}{\mathbf{D}}\|^{2}_{F}.$ (3.17)
@@
$\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}$ (3.7)${\leq}\|P_{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{1}\left(\|P_{\mathbf{A}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\right)+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
(3.4)${\leq}c_{1}\left(\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+\lambda\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}_{0}{\mathbf{D}}\|^{2}_{F}\right)+2\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq c_{1}\|P_{\mathbf{A}}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+2\lambda c_{1}\|P_{\mathbf{A}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\overset{c_{2}=\max(2\lambda,1)}{\leq}c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq 2c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+2c_{1}c_{2}\|{\mathbf{X}}_{0}\|^{2}_{F}\|{\mathbf{I}}-{\mathbf{D}}\|^{2}+(2\lambda c_{1}+2)\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq 2c_{1}c_{2}(1+2\varepsilon^{2})\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+4c_{1}c_{2}\varepsilon^{2}\|{\mathbf{X}}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{\mathbf{A}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
@@ 4.1. General convergence for sketched multiplicative updates
Multiplicative updates has been one of the most popular algorithms for NMF since the introduction in [22]. In this section we show how to modify the classical multiplicative updates algorithm for the various objectives we have derived in earlier sections.
To this end, we prove a general theorem for multiplicative updates for multiple minimization terms. We will see in the next Section 4.2 that giving nonnegativity conditions on the sums (4.2) (rather than the stronger conditions on the individual terms) matters so we can put realistic assumption on our regularization terms that include orthogonal projection operators.
Theorem 4.1.
@@
$\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\frac{1}{2}\sum_{i=1}^{s}\|{\mathbf{A}}_{i}({\mathbf{X}}^{({\mathbf{A}})}_{i}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}+\frac{1}{2}\sum_{j=1}^{t}\|({\mathbf{X}}^{({\mathbf{B}})}_{j}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{B}}_{j}\|_{F}^{2},$ (4.1)
where all ${\mathbf{X}}_{i}^{({\mathbf{A}})},{\mathbf{X}}_{j}^{({\mathbf{B}})}\in{\mathbb{R}}^{m\times n}$, ${\mathbf{A}}_{i}\in{\mathbb{R}}^{k\times m}$, ${\mathbf{B}}_{j}\in{\mathbb{R}}^{n\times k}$ and $({\mathbf{U}},{\mathbf{V}}\geq 0)$ means $({\mathbf{U}}\in{\mathbb{R}}_{+}^{m\times r},{\mathbf{V}}\in{\mathbb{R}}_{+}^{n\times r}).$ Let ${\mathbf{U}}$ and ${\mathbf{V}}$ be $m\times r$ and $n\times r$ matrices respectively. If all six matrices
$\left\{{\mathbf{U}},{\mathbf{V}},(\sum_{i=1}^{s}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}),(\sum_{j=1}^{t}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}),(\sum_{i=1}^{s}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{X}}^{({\mathbf{A}})}_{i}),(\sum_{j=1}^{t}{\mathbf{X}}^{({\mathbf{B}})}_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T})\right\}$ (4.2)
are entry-wise nonnegative, then the objective (4.1) is nonincreasing under the updates
@@
Remark 4.2.
(Implementation considerations) Note that we can compute the matrix ${\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}$ by the multiplication order given by ${\mathbf{A}}_{i}^{T}(({\mathbf{A}}_{i}{\mathbf{U}})({\mathbf{V}}^{T}{\mathbf{V}}))$. This order never requires us to store a matrix of size larger than $\max(n,m)\times\max(r,k)$. Similar procedures can be used for other terms appearing in the updates of Theorem 4.1.
The proof of the theorem follows the standard approach that justifies the validity of multiplicative updates algorithm for NMF problem (e.g., [22]), where nonnegativity assumption ensures its validity on the sketched problem.
Proof of Theorem 4.1.
@@
Using the mixed Kronecker matrix-vector product property $({\mathbf{M}}_{1}\otimes{\mathbf{M}}_{2})\text{vec}({\mathbf{U}})=\text{vec}({\mathbf{M}}_{2}{\mathbf{U}}{\mathbf{M}}_{1}^{T})$ that holds for any appropriately sized matrices ${\mathbf{U}},{\mathbf{M}}_{1},{\mathbf{M}}_{2}$, we can rewrite the objective function (4.1) as
@@
$G({\mathbf{u}}^{\prime},{\mathbf{u}})=F({\mathbf{u}})+({\mathbf{u}}^{\prime}-{\mathbf{u}})^{T}\nabla F({\mathbf{u}})+\frac{1}{2}({\mathbf{u}}^{\prime}-{\mathbf{u}})^{T}{\mathbf{K}}_{\mathbf{u}}({\mathbf{u}}^{\prime}-{\mathbf{u}}),$ (4.3)
where the matrix ${\mathbf{K}}_{\mathbf{u}}$ is a diagonal matrix with the diagonal $({\mathbf{W}}^{T}{\mathbf{W}}{\mathbf{u}})/{\mathbf{u}}$, recall that /// represents elementwise division. We claim that $G$ majorizes $F$, i.e. $G({\mathbf{u}},{\mathbf{u}})=F({\mathbf{u}})$ and $G({\mathbf{u}}^{\prime},{\mathbf{u}})\geq F({\mathbf{u}}^{\prime})$. It is clear that $G({\mathbf{u}},{\mathbf{u}})=F({\mathbf{u}})$. We can write
@@
from the comparison of (4.3) with the Taylor decomposition for the quadratic function $F({\mathbf{u}}^{\prime})$ at ${\mathbf{u}}$. So, to check that $G({\mathbf{u}}^{\prime},{\mathbf{u}})\geq F({\mathbf{u}}^{\prime})$, it is sufficient to show that the matrix ${\mathbf{K}}_{\mathbf{u}}-{\mathbf{W}}^{T}{\mathbf{W}}$ is positive semidefinite. Equivalently, it is sufficient to show that
@@
is positive semidefinite. Indeed, for any vector ${\mathbf{z}}$ of the appropiate size, ${\mathbf{z}}^{T}{\mathbf{M}}{\mathbf{z}}=({\mathbf{z}}\circ{\mathbf{u}})^{T}({\mathbf{K}}_{\mathbf{u}}-{\mathbf{W}}^{T}{\mathbf{W}})({\mathbf{z}}\circ{\mathbf{u}})$, recall that $\circ$ defines element-wise product.
@@
$\arg\min_{{\mathbf{u}}^{\prime}}G({\mathbf{u}}^{\prime},{\mathbf{u}})={\mathbf{u}}-{\mathbf{K}}_{\mathbf{u}}^{-1}\nabla F({\mathbf{u}})={\mathbf{u}}\circ\frac{{\mathbf{W}}^{T}{\mathbf{y}}}{{\mathbf{W}}^{T}{\mathbf{W}}{\mathbf{u}}}.$ (4.4)

### v2

@@ 3.4. Nonnegativity in compression
In the further sections, we consider the variations of multiplicative updates algorithm to iteratively minimize our proposed sketched objective functions. The multiplicative updates algorithm is valid due to the fact that the matrices involved in the iterative process are nonnegative. Note that this convenient property is destroyed by sketching unless we have that ${\mathbf{A}}^{T}{\mathbf{A}}$ is an element-wise nonnegative matrix. However, the nonnegativity of the sketching matrix is not expected to hold neither for approximately orthonormal random sketches nor for the data-adapted sketching matrices coming from randomized rangefinder algorithm.
In this section, we show how to overcome this issue: it actually suffices to add some extra penalty terms taking the form
(16) $\sigma\|\mathbf{1}_{m}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2}\quad\text{and/or}\quad\sigma\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\mathbf{1}_{n}\|^{2},$
@@
Corollary 2.
Suppose $\mathbf{X}$} has an exact nonnegative factorization ${\mathbf{X}}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$, where ${\mathbf{U}}_{0}\in{\mathbb{R}}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}^{n\times r}$ with $r\leq\min\{n,m\}$. Let ${\mathbf{A}}_{1}$ and ${\mathbf{A}}_{2}$ be generic random matrices of sizes $k_{1}\times m$ and $n\times k_{2}$, respectively, with $k_{1},k_{2}\geq r$. If for some $\lambda_{1},\lambda_{2}>0$ and $\sigma_{1},\sigma_{2}\geq 0$
(17) $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}{\color[rgb]{0,0,0}\;\in\;}\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}L$ $({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})$
@@
where $L({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}):=\|{\mathbf{A}}_{1}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}+\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{A}}_{2}\|_{F}^{2}+\lambda_{1}\|P_{{\mathbf{X}}{\mathbf{A}}_{2}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|_{F}^{2}+\lambda_{2}\|{\mathbf{U}}{\mathbf{V}}^{T}P_{{\mathbf{X}}^{T}{\mathbf{A}}_{1}^{T}}^{\perp}\|_{F}^{2},$ then ${\mathbf{X}}=\tilde{{\mathbf{U}}}\tilde{{\mathbf{V}}}^{T}$.
@@
and similarly for adding the term $\sigma_{2}\|({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\mathbf{1}_{n}\|^{2}_{F}$. Then the statement follows directly from Theorem 1. ∎
When we do not assume that $\mathbf{X}$} has an exact nonnegative decomposition of dimension $r$, adding the regularizer of the form (16) is still possible under an additional condition, essentially imposing that the column-sums of $\mathbf{X}$} and ${\mathbf{U}}{\mathbf{V}}^{T}$ approximately match if ${\mathbf{U}},{\mathbf{V}}$ are optimal to (1).
Corollary 3.
Let ${\mathbf{X}}\in{\mathbb{R}}_{+}^{m\times n}$ be a nonnegative matrix and ${\mathbf{A}}\in{\mathbb{R}}^{k\times m}$ is a matrix with orthogonal rows. Let ${\mathbf{U}}_{0}\in{\mathbb{R}}_{+}^{m\times r}$, ${\mathbf{V}}_{0}\in{\mathbb{R}}_{+}^{r\times n}$ give a solution to the original NMF problem (1) and ${\mathbf{X}}_{0}={\mathbf{U}}_{0}{\mathbf{V}}_{0}^{T}$. Additionally assume that $\left\|\mathbf{1}^{T}({\mathbf{X}}_{0}-{\mathbf{X}})/\mathbf{1}^{T}{\mathbf{X}}\right\|\leq\varepsilon<1$, where // denotes element-wise division and $\mathbf{1}\in{\mathbb{R}}^{m}$ is the vector of all ones. If $\lambda,\sigma>0,$
(18) $\tilde{{\mathbf{U}}},\tilde{{\mathbf{V}}}{\color[rgb]{0,0,0}\;\in\;}\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\left\|{\mathbf{A}}\left({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T}\right)\right\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{U}}{\mathbf{V}}^{T}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-{\mathbf{U}}{\mathbf{V}}^{T})\|^{2},$
@@
$\frac{\|{\mathbf{X}}-\tilde{{\mathbf{X}}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}\leq c_{\lambda}\left[\frac{\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\frac{\|{\mathbf{X}}-P_{{\mathbf{A}}^{T}}{\mathbf{X}}\|^{2}_{F}}{\|{\mathbf{X}}\|^{2}_{F}}+\varepsilon^{2}\right].$
For example, if we have that $\lambda,\varepsilon\in(0,1/2)$ one can bound ${\color[rgb]{0,0,0}c_{\lambda}}=\max\{6,8/\lambda\}.$
@@
Let ${\mathbf{D}}\in{\mathbb{R}}^{n\times n}$ be the diagonal matrix with nonzero entries given by $(\mathbf{1}^{T}{\mathbf{X}}/\mathbf{1}^{T}{\mathbf{X}}_{0})$, and so $\mathbf{1}^{T}{\mathbf{X}}_{0}{\mathbf{D}}=\mathbf{1}^{T}{\mathbf{X}}$. Note that ${\mathbf{X}}_{0}{\mathbf{D}}$ is some (not necessarily optimal) solution to (18) and so we have
$\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}$ $\leq\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}+\sigma\|\mathbf{1}^{T}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}$
(19) $\leq\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}_{0}{\mathbf{D}}\|^{2}_{F}.$
@@
[…]
$\leq c_{1}\left(\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-\tilde{{\mathbf{X}}})\|^{2}_{F}+\lambda\|P_{{\mathbf{A}}^{T}}^{\perp}\tilde{{\mathbf{X}}}\|^{2}_{F}\right)+2\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
[…]
$\leq c_{1}\|P_{{\mathbf{A}}^{T}}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+2\lambda c_{1}\|P_{{\mathbf{A}}^{T}}^{\perp}({\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}})\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\overset{c_{2}=\max(2\lambda,1)}{\leq}c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}{\mathbf{D}}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq 2c_{1}c_{2}\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+2c_{1}c_{2}\|{\mathbf{X}}_{0}\|^{2}_{F}\|{\mathbf{I}}-{\mathbf{D}}\|^{2}+(2\lambda c_{1}+2)\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}$
$\leq 2c_{1}c_{2}(1+2\varepsilon^{2})\|{\mathbf{X}}-{\mathbf{X}}_{0}\|^{2}_{F}+4c_{1}c_{2}\varepsilon^{2}\|{\mathbf{X}}\|^{2}_{F}+(2\lambda c_{1}+2)\|P_{{\mathbf{A}}^{T}}^{\perp}{\mathbf{X}}\|^{2}_{F}.$
This completes the proof of Corollary 3. ∎
Remark 5.
We note that the statement of Corollary 3 holds with any $k,r\geq 1$. However, in general, $k$ needs to upper-bound $r$ for the error terms to be controllable, in the same way as discussed in Theorem 3 above.
@@ 4.1. General convergence for sketched multiplicative updates
Multiplicative updates (MU) has been one of the most popular algorithms for the standard NMF problem (1) since the introduction in [25]. In this section we show how to modify the classical MU algorithm for the various objectives we have derived in earlier sections.
To this end, we first prove a general Theorem 5 for MU for multiple minimization terms (instead of just one as in [25]) after possible additional linear transformations on the left and on the right (20). Then – in Corollaries 4, 5, 6 – we demonstrate several special cases of this generalized multiplicative updates scheme that extend the applicability of the MU framework specifically to our proposed compressed NMF problems, as additional linear transformations allow to represent projection and shift terms.
Theorem 5.
@@
(20) $\operatorname*{arg\,min}_{{\mathbf{U}},{\mathbf{V}}\geq 0}\frac{1}{2}\sum_{i=1}^{s}\|{\mathbf{A}}_{i}({\mathbf{X}}^{({\mathbf{A}})}_{i}-{\mathbf{U}}{\mathbf{V}}^{T})\|_{F}^{2}+\frac{1}{2}\sum_{j=1}^{t}\|({\mathbf{X}}^{({\mathbf{B}})}_{j}-{\mathbf{U}}{\mathbf{V}}^{T}){\mathbf{B}}_{j}\|_{F}^{2},$
where all ${\mathbf{X}}_{i}^{({\mathbf{A}})},{\mathbf{X}}_{j}^{({\mathbf{B}})}\in{\mathbb{R}}^{m\times n}$, ${\mathbf{A}}_{i}\in{\mathbb{R}}^{k\times m}$, ${\mathbf{B}}_{j}\in{\mathbb{R}}^{n\times k}$ and $({\mathbf{U}},{\mathbf{V}}\geq 0)$ means $({\mathbf{U}}\in{\mathbb{R}}_{+}^{m\times r},{\mathbf{V}}\in{\mathbb{R}}_{+}^{n\times r}).$ If all six matrices
(21) $\left\{{\mathbf{U}},{\mathbf{V}},(\sum_{i=1}^{s}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}),(\sum_{j=1}^{t}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T}),(\sum_{i=1}^{s}{\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{X}}^{({\mathbf{A}})}_{i}),(\sum_{j=1}^{t}{\mathbf{X}}^{({\mathbf{B}})}_{j}{\mathbf{B}}_{j}{\mathbf{B}}_{j}^{T})\right\}$
are entry-wise nonnegative, then the objective (20) is nonincreasing under the updates
@@
Remark 6.
(Implementation considerations) Note that we can compute the matrix ${\mathbf{A}}_{i}^{T}{\mathbf{A}}_{i}{\mathbf{U}}{\mathbf{V}}^{T}{\mathbf{V}}$ by the multiplication order given by ${\mathbf{A}}_{i}^{T}(({\mathbf{A}}_{i}{\mathbf{U}})({\mathbf{V}}^{T}{\mathbf{V}}))$. This order never requires us to store a matrix of size larger than $\max(n,m)\times\max(r,k)$. Similar procedures can be used for other terms appearing in the updates of Theorem 5.
The proof of the theorem follows the standard approach that justifies the validity of multiplicative updates algorithm for NMF problem (e.g., [25]), where nonnegativity assumption ensures its validity on the sketched problem. We remark that giving nonnegativity conditions on the sums (21) (rather than the stronger conditions on the individual terms) matters so we can put realistic assumption on our regularization terms that include orthogonal projection operators (in Section 4.2).
Proof of Theorem 5.
@@
Using the mixed Kronecker matrix-vector product property $({\mathbf{M}}_{1}\otimes{\mathbf{M}}_{2})\text{vec}({\mathbf{U}})=\text{vec}({\mathbf{M}}_{2}{\mathbf{U}}{\mathbf{M}}_{1}^{T})$ that holds for any appropriately sized matrices ${\mathbf{U}},{\mathbf{M}}_{1},{\mathbf{M}}_{2}$, we can rewrite the objective function (20) as
@@
(22) $G({\mathbf{u}}^{\prime},{\mathbf{u}})=F({\mathbf{u}})+({\mathbf{u}}^{\prime}-{\mathbf{u}})^{T}\nabla F({\mathbf{u}})+\frac{1}{2}({\mathbf{u}}^{\prime}-{\mathbf{u}})^{T}{\mathbf{K}}_{\mathbf{u}}({\mathbf{u}}^{\prime}-{\mathbf{u}}),$
where the matrix ${\mathbf{K}}_{\mathbf{u}}$ is a diagonal matrix with the diagonal $({\mathbf{W}}^{T}{\mathbf{W}}{\mathbf{u}})/{\mathbf{u}}$, recall that // represents elementwise division. We claim that $G$ majorizes $F$, i.e., $G({\mathbf{u}},{\mathbf{u}})=F({\mathbf{u}})$ and $G({\mathbf{u}}^{\prime},{\mathbf{u}})\geq F({\mathbf{u}}^{\prime})$. It is clear that $G({\mathbf{u}},{\mathbf{u}})=F({\mathbf{u}})$. We can write
@@
from the comparison of (22) with the Taylor decomposition for the quadratic function $F({\mathbf{u}}^{\prime})$ at $\mathbf{u}$}. So, to check that $G({\mathbf{u}}^{\prime},{\mathbf{u}})\geq F({\mathbf{u}}^{\prime})$, it is sufficient to show that the matrix ${\mathbf{K}}_{\mathbf{u}}-{\mathbf{W}}^{T}{\mathbf{W}}$ is positive semidefinite. Equivalently, it is sufficient to show that
@@
is positive semidefinite. Indeed, for any vector $\mathbf{z}$} of the appropriate size, ${\mathbf{z}}^{T}{\mathbf{M}}{\mathbf{z}}=({\mathbf{z}}\circ{\mathbf{u}})^{T}({\mathbf{K}}_{\mathbf{u}}-{\mathbf{W}}^{T}{\mathbf{W}})({\mathbf{z}}\circ{\mathbf{u}})$, recall that $\circ$ defines element-wise product.
@@
(23) $\arg\min_{{\mathbf{u}}^{\prime}}G({\mathbf{u}}^{\prime},{\mathbf{u}})={\mathbf{u}}-{\mathbf{K}}_{\mathbf{u}}^{-1}\nabla F({\mathbf{u}})={\mathbf{u}}\circ\frac{{\mathbf{W}}^{T}{\mathbf{y}}}{{\mathbf{W}}^{T}{\mathbf{W}}{\mathbf{u}}}.$